A string is pulled under tension of 50 newton. A transverse wave setup on the string can be represented by y=0.25 sin (2.28x - 966).Calculate the mass per unit length of the string.
1
Expert's answer
2018-10-25T15:48:09-0400

The speed of the wave:

v=√(T/μ)

v=966/2.28 m/s.

The mass per unit length of the string is

μ=F/v^2 =50/(966/2.28)^2 = 0.000279 kg/m=0.279 g/m